# quantum thermodynamics, thermal operations, thermomajorization, thermal equilibrium, Gibbs-stochastic matrix, cyclic process

### v5

@@ Abstract
Drawing inspiration from transportation theory, in this work we introduce the notions of “well-structured” and “stable” Gibbs states and we investigate their implications for quantum thermodynamics and its resource theory approach via thermal operations. It turns out that, in the quasi-classical realm, global cyclic state transfers are impossible if and only if the Gibbs state is stable. Moreover, using a geometric approach by studying the so-called thermomajorization polytope we prove that any subspace in equilibrium can be brought out of equilibrium via thermal operations. Interestingly, the case of some subsystem being in equilibrium can be witnessed via degenerate extreme points of the thermomajorization polytope, assuming the Gibbs state of the system is well structured. These physical considerations are complemented by simple new constructions for the polytope’s extreme points as well as for an important class of extremal Gibbs-stochastic matrices.
@@ 2.1 Thermomajorization for Quasi-Classical States
$\begin{split}\mathrm{th}_{d,y}(c)&=\min_{\{i=1,\ldots,n\,:\,d_{i}>0\}}\Big{(}\Big{(}\sum_{j=1}^{n}\max\Big{\{}y_{j}-\frac{y_{i}}{d_{i}}d_{j},0\Big{\}}\Big{)}+\frac{y_{i}}{d_{i}}c\Big{)}\\ &=\frac{{\bf e}^{\top}y}{2}+\min_{\{i=1,\ldots,n\,:\,d_{i}>0\}}\Big{(}\Big{\|}y-\frac{y_{i}}{d_{i}}d\Big{\|}_{1}+\frac{y_{i}}{d_{i}}\Big{(}c-\frac{{\bf e}^{\top}d}{2}\Big{)}\Big{)}\end{split}$ (2)
where, here and henceforth, ${\bf e}:=(1,\ldots,1)^{\top}$ and $\|\cdot\|_{1}$ is the usual vector $1$-norm. Note that this curve is a generalization of the notion of Lorenz curves from majorization theory (Marshall et al., 2011, p. 5). While the condition $d_{i}>0$ from the minimum in (2) is redundant for now, it allows us to generalize the definition of $\mathrm{th}_{d,y}$ to the zero-temperature case where some of the entries of $d$ vanish, cf. Remark 2.1 below. {Remark} It is clear from the definition that the thermomajorization curve is invariant under permutations in the sense that $\mathrm{th}_{\underline{\sigma}d,\underline{\sigma}y}\equiv\mathrm{th}_{d,y}$ for all $\sigma\in S_{n}$. Here, given some permutation ${\sigma}\in S_{n}$ we write $\underline{\sigma}$ for the corresponding permutation matrix $\sum_{i=1}^{n}e_{i}e_{{\sigma}(i)}^{\top}$. In particular the identities $\underline{{\sigma}\circ\tau}=\underline{\tau}\cdot\underline{{\sigma}}$, $(\underline{{\sigma}}x)_{j}=x_{{\sigma}(j)}$, and $(\underline{\sigma}A\underline{\tau})_{jk}=A_{\sigma(j)\tau^{-1}(k)}$ hold for all $A\in\mathbb{R}^{n\times n}$, $x\in\mathbb{R}^{n}$, $j,k=1,\ldots,n$, and all $\sigma,\tau\in S_{n}$. Now the precise connection between thermomajorization and (quasi-classical) state transfers is summarized in the following well-known result: Given any $x,y\in\mathbb{R}^{n}$ the following statements are equivalent (vom Ende and Dirr, 2022, Prop. 1).
@@
    $\sum_{i=1}^{j}x_{\tau(i)}\leq\min_{\{i=1,\ldots,n\,:\,d_{i}>0\}}\Big{(}\Big{(}\sum_{j=1}^{n}\max\Big{\{}y_{j}-\frac{y_{i}}{d_{i}}d_{j},0\Big{\}}\Big{)}+\frac{y_{i}}{d_{i}}\Big{(}\sum_{k=1}^{j}d_{\tau(k)}\Big{)}\Big{)}$
@@
[Thermomajorization for Zero Temperature] The above conditions for thermomajorization remain well defined even if some entries of $d$ vanish, assuming $\mathrm{th}_{d,y}$ is defined via (2). Physically, this relates to the temperature being zero in which case all entries of $d$ that do not correspond to the zero-point energy of the system vanish. If this lowest energy level with corresponding index $j$ is non-degenerate, then $d=d(T)\to e_{j}$ as $T\to 0^{+}$. Interestingly, the above characterization of thermomajorization stays valid in this regime. For a proof we refer to App. A. This explains known continuity problems of the associated polytope (cf. Example 3 in vom Ende and Dirr (2022)): If some of the $d_{i}$ vanish, so do the corresponding inequalities meaning there are less restrictions. Hence the polytope can only become larger in this case. The importance of the $1$-norm conditions is that they lead to a beautiful characterization of $M_{d}(y)$ from (1), cf. (vom Ende and Dirr, 2022, Thm. 10):
@@ 3.1 Cyclic State Transfers and Subspaces in Equilibrium
Because $\mathrm{th}_{d,y}$ is concave and because $m^{\top}d\in(0,{\bf e}^{\top}d)$ by the assumptions on $m$ and $d$, Lemma B (iv) (App. B) shows that (6) can only hold if $\mathrm{th}_{d,y}$ is linear. But the latter is equivalent to $y$ being a multiple of $d$ as, by definition, the slopes of $\mathrm{th}_{d,y}$ are given by ${y_{j}}/{d_{j}}$.
@@ 3.2 Extreme Points of the Thermomajorization Polytope
Assume $\sigma\in S_{n}$ satisfies $E_{d,y}(\sigma)=y$ so $(E_{d,y}(\sigma))_{\sigma(j)}=y_{\sigma(j)}$ for all $j=1,\ldots,n$. By definition of $E_{d,y}$ this means $\mathrm{th}_{d,y}(\sum_{i=1}^{j}d_{\sigma(i)})=\mathrm{th}_{d,y}(\sum_{i=1}^{j-1}d_{\sigma(i)})+y_{\sigma(j)}$ which by induction is equivalent to $\mathrm{th}_{d,y}(\sum_{i=1}^{j}d_{\sigma(i)})=\sum_{i=1}^{j}y_{\sigma(i)}$ for all $j=1,\ldots,n-1$. But by Lemma B (App. B) this holds if and only if ${y_{\sigma(1)}}/{d_{\sigma(1)}}\geq\ldots\geq{y_{\sigma(n)}}/{d_{\sigma(n)}}$. ∎
@@ 5 Conclusions
We acknowledge support by the Open Access Publication Fund of Freie Universität Berlin.
[…]
@@ Acknowledgements.
Appendix A Thermomajorization in the Zero-Temperature Limit
Before proving the result in question we first need the following lemma. {Lemma} Let $x,y,d\in\mathbb{R}^{n}$ with $d>0$ as well as numbers $x_{+},x_{-},y_{+},y_{-}\geq 0$ be given. If ${\bf e}^{\top}x+x_{+}-x_{-}={\bf e}^{\top}y+y_{+}-y_{-}$ and
$\|d_{i}x-y_{i}d\|_{1}+d_{i}(x_{+}+x_{-})\leq\|d_{i}y-y_{i}d\|_{1}+d_{i}(y_{+}+y_{-})$ (11)
for all $i=1,\ldots,n$, then $x_{-}\leq y_{-}$ and $x_{+}\leq y_{+}$.
Proof.
Combining (11) with the trace condition yields
$\|d_{i}x-y_{i}d\|_{1}+d_{i}(x_{+}+x_{-})\leq\,\,$ $\|d_{i}y-y_{i}d\|_{1}+d_{i}\big{(}y_{+}+({\bf e}^{\top}y+y_{+}-{\bf e}^{\top}x-x_{+}+x_{-})\big{)}$
which is equivalent to $\|d_{i}x-y_{i}d\|_{1}\leq\|d_{i}y-y_{i}d\|_{1}+d_{i}{\bf e}^{\top}(y-x)+2d_{i}(y_{+}-x_{+})$ and thus
$y_{+}-x_{+}\geq\frac{\|d_{i}x-y_{i}d\|_{1}-\|d_{i}y-y_{i}d\|_{1}-d_{i}{\bf e}^{\top}(y-x)}{2d_{i}}\,.$ (12)
Hence all we have to find is $i\in\{1,\ldots,n\}$ such that the r.h.s. of (12) is non-negative. For this let $i=1,\ldots,n$ be the index which satisfies $\frac{y_{i}}{d_{i}}=\max_{j=1,\ldots,n}\frac{y_{j}}{d_{j}}$. Then one finds $d_{i}y_{j}-y_{i}d_{j}=d_{i}d_{j}(\frac{y_{j}}{d_{j}}-\frac{y_{i}}{d_{i}})\leq 0$ for all $j=1,\ldots,n$, hence $\|d_{i}y-y_{i}d\|_{1}=-{\bf e}^{\top}(d_{i}y-y_{i}d)$. Therefore the numerator of the r.h.s. of (12) for this index $i$ is equal to $\|d_{i}x-y_{i}d\|_{1}+{\bf e}^{\top}(d_{i}y-y_{i}d)-{\bf e}^{\top}(d_{i}y)+{\bf e}^{\top}(d_{i}x)=\|d_{i}x-y_{i}d\|_{1}+{\bf e}^{\top}(d_{i}x-y_{i}d)\geq 0$. The inequality $x_{-}\leq y_{-}$ is shown analogously by replacing $y_{+}$ (instead of $y_{-}$) in (11) via the trace condition. One then shows $\|d_{i}x-y_{i}d\|_{1}-\|d_{i}y-y_{i}d\|_{1}+d_{i}{\bf e}^{\top}(y-x)\geq 0$ similar to our argument above by choosing $i$ to be the index such that $\frac{y_{i}}{d_{i}}=\min_{j=1,\ldots,n}\frac{y_{j}}{d_{j}}$ because then $\|d_{i}y-y_{i}d\|_{1}={\bf e}^{\top}(d_{i}y-y_{i}d)$. ∎
With this we can show that the characterizations of thermomajorization for quasi-classical states (Sec. 2.1) extends to the case of temperature zero, assuming that the lowest energy level of the system is non-degenerate. What this means mathematically is that the (normalized) Gibbs vector becomes some standard basis vector $e_{j}$ in the limit $T\to 0^{+}$. W.l.o.g. $j=1$ because given any permutation matrix $\underline{\sigma}\in\mathbb{R}^{n\times n}$, $x\prec_{d}y$ is equivalent $\underline{\sigma}x\prec_{\underline{\sigma}d}\underline{\sigma}y$. We claim that the following statements are equivalent for all $x,y\in\mathbb{R}^{n}$:
  • (i)
    There exists $A\in\mathbb{R}_{+}^{n\times n}$ such that ${\bf e}^{\top}A={\bf e}^{\top}$, $Ae_{1}=e_{1}$, and $Ay=x$.
  • (ii)
    ${\bf e}^{\top}x={\bf e}^{\top}y$, and for all $t\in\mathbb{R}$ one has $\|x-te_{1}\|_{1}\leq\|y-te_{1}\|_{1}$.
  • (iii)
    ${\bf e}^{\top}x={\bf e}^{\top}y$ and $\|x-y_{1}e_{1}\|_{1}\leq\|y-y_{1}e_{1}\|_{1}$.
  • (iv)
    ${\bf e}^{\top}x={\bf e}^{\top}y$, and for all $c\in[0,{\bf e}^{\top}d]$ one has $\mathrm{th}_{d,x}(c)\leq\mathrm{th}_{d,y}(c)$.
Proof.
“(ii) $\Leftrightarrow$ (iv)”: Eq (3). “(i) $\Rightarrow$ (ii)”: By assumption, ${\bf e}^{\top}x=({\bf e}^{\top}A)y={\bf e}^{\top}y$ and $x-y_{j}e_{j}=Ay-y_{j}Ae_{j}=A(y-y_{j}e_{j})\,.$ This together with the fact that every column-stochastic matrix is $1$-norm contractive (vom Ende and Dirr, 2022, Remark 1 (i)) implies (ii). “(ii) $\Rightarrow$ (iii)”: Trivial. “(iii) $\Rightarrow$ (i)”: The idea is as follows: partition $y=\scriptsize\begin{pmatrix}y_{1}\\ y^{\prime}\end{pmatrix}$, $x=\scriptsize\begin{pmatrix}x_{1}\\ x^{\prime}\end{pmatrix}$ where $x^{\prime},y^{\prime}\in\mathbb{R}^{n-1}$, and perform the following chain of transformations:
[…]
Before explaining the undefined objects let us make the following key observation: If some entries of $d$ are zero then any $d$-stochastic matrix has to be block triangular. More precisely given $A\in\mathbb{R}_{+}^{n\times n}$ and $d=\scriptsize\begin{pmatrix}\tilde{d}\\ 0_{n-m}\end{pmatrix}$ where $\tilde{d}\in\mathbb{R}^{m}$, $\tilde{d}>0$, the matrix $A$ is $d$-stochastic if and only if there exist a $\tilde{d}$-stochastic matrix $\tilde{A}\in\mathbb{R}_{+}^{m\times m}$ and probability vectors $v_{1},\ldots,v_{k}\in\mathbb{R}_{+}^{n}$ (i.e. ${\bf e}^{\top}v_{1}=\ldots={\bf e}^{\top}v_{k}=1$) such that
$A=\left(\begin{matrix}\tilde{A}\\ 0_{(n-m)\times m}\end{matrix}\ \;\begin{matrix}v_{1}&\cdots&v_{k}\end{matrix}\right)\,.$
Now let us go over the above instructions step by step:
  • (1)
    Define row vectors $w_{+},w_{-}\in\mathbb{R}^{n-1}$ as follows: given $i=1,\ldots,n-1$ set $(w_{+})_{i}:=1$ if $y^{\prime}_{i}\geq 0$, else $(w_{+})_{i}:=0$. Then set $w_{-}:={\bf e}^{\top}-w_{+}$. Thus $w_{+}$ indicates the location of the non-negative entries of $y^{\prime}$, and $w_{-}$ tells us where the negative entries of $y^{\prime}$ are. Now the column-stochastic matrix
    $A_{1}:=\left(\begin{matrix}1\\ 0\\ 0\\ 0\end{matrix}\ \;\begin{matrix}0\\ w_{+}\\ w_{-}\\ 0\end{matrix}\right)$
    collects the positive entries of $y^{\prime}$ in the second position, and the negative entries of $y^{\prime}$ in the third position131313 We may w.l.o.g. assume $n\geq 3$ because we can append as many zeros to the original vectors as we want without altering the problem: one readily verifies that $x\prec_{d}y$ is equivalent to $\scriptsize\begin{pmatrix}x\\ 0_{k}\end{pmatrix}\prec_{\begin{pmatrix}d\\ 0_{k}\end{pmatrix}}\begin{pmatrix}y\\ 0_{k}\end{pmatrix}$ for all $x,y,d\in\mathbb{R}^{n}$, $d\geq 0$, and $k\in\mathbb{N}$.. We write $y_{+}:=w_{+}y^{\prime}$ and $y_{-}:=-(w_{-}y^{\prime})$.
  • (2)
    Similarly, define $x_{+}:=\sum_{i=1}^{k}\max\{x_{i}^{\prime},0\}$, $x_{-}:=-\sum_{i=1}^{k}\min\{x_{i}^{\prime},0\}$. Note $x_{\pm},y_{\pm}\geq 0$ so evaluating the $1$-norm inequality from (iii) readily implies the assumptions of Lemma A (when setting $n=1$, $x:=x_{1}$, $y:=y_{1}$, $d:=1$). Hence Lemma A shows $x_{-}\leq y_{-}$, $x_{+}\leq y_{+}$. This implies that
    $A_{2}:=\begin{pmatrix}1&1-\frac{x_{+}}{y_{+}}&1-\frac{x_{-}}{y_{-}}\\ 0&\frac{x_{+}}{y_{+}}&0\\ 0&0&\frac{x_{-}}{y_{-}}\\ 0&0&0\end{pmatrix}$
    has non-negative entries so $A_{2}$ is column-stochastic, has $e_{j}$ as fixed point, and carries out the desired transformation. Note that if $y_{+}$ ($y_{-}$) is zero, so is $x_{+}$ ($x_{-}$) meaning we can set $\frac{x_{+}}{y_{+}}:=1$ ($\frac{x_{-}}{y_{-}}:=1$). Concluding this step, ${\bf e}^{\top}x={\bf e}^{\top}y$ implies $y_{1}+(y_{+}-x_{+})-(y_{-}-x_{-})=x_{1}$.
  • (3)
    Similarly to (1) define $A_{3}$ such that the bottom right $(n-1)\times(n-1)$ block of $A_{3}$ redistributes the accumulated positive and negative entries of $x^{\prime}$ back to its original place.∎
Based on this result it seems reasonable to conjecture that for all $d\in\mathbb{R}_{+}^{n}$, $x\prec_{d}y$ is equivalent to ${\bf e}^{\top}x={\bf e}^{\top}y$ and $\|d_{i}x-y_{i}d\|_{1}\leq\|d_{i}y-y_{i}d\|_{1}$, $i=1,\ldots,n$. However, proving this seems to not be as straightforward as the above non-degenerate temperature-zero case, and it would be beside the point of this article either way.
Appendix B Basic Properties of Concave Functions

### v6

@@ Abstract
Drawing inspiration from transportation theory, in this work we introduce the notions of “well-structured” and “stable” Gibbs states and we investigate their implications for quantum thermodynamics and its resource theory approach via thermal operations. It turns out that, in the quasi-classical realm, global cyclic state transfers are impossible if and only if the Gibbs state is stable. Moreover, using a geometric approach by studying the so-called thermomajorization polytope we prove that any subsystem in equilibrium can be brought out of equilibrium via thermal operations. Interestingly, the case of some subsystem being in equilibrium can be witnessed via degenerate extreme points of the thermomajorization polytope, assuming the Gibbs state of the system is well structured. These physical considerations are complemented by simple new constructions for the polytope’s extreme points as well as for an important class of extremal Gibbs-stochastic matrices.
@@ 2.1 Thermomajorization for Quasi-Classical States
$\begin{split}\mathrm{th}_{d,y}(c)&=\min_{\{i=1,\ldots,n\}}\Big{(}\Big{(}\sum_{j=1}^{n}\max\Big{\{}y_{j}-\frac{y_{i}}{d_{i}}d_{j},0\Big{\}}\Big{)}+\frac{y_{i}}{d_{i}}c\Big{)}\\ &=\frac{{\bf e}^{\top}y}{2}+\min_{\{i=1,\ldots,n\}}\Big{(}\Big{\|}y-\frac{y_{i}}{d_{i}}d\Big{\|}_{1}+\frac{y_{i}}{d_{i}}\Big{(}c-\frac{{\bf e}^{\top}d}{2}\Big{)}\Big{)}\end{split}$ (2)
where, here and henceforth, ${\bf e}:=(1,\ldots,1)^{\top}$ and $\|\cdot\|_{1}$ is the usual vector $1$-norm. Note that this curve is a generalization of the notion of Lorenz curves from majorization theory (Marshall et al., 2011, p. 5).
{Remark}
It is clear from the definition that the thermomajorization curve is invariant under permutations in the sense that $\mathrm{th}_{\underline{\sigma}d,\underline{\sigma}y}\equiv\mathrm{th}_{d,y}$ for all $\sigma\in S_{n}$. Here, given some permutation ${\sigma}\in S_{n}$ we write $\underline{\sigma}$ for the corresponding permutation matrix $\sum_{i=1}^{n}e_{i}e_{{\sigma}(i)}^{\top}$. In particular the identities $\underline{{\sigma}\circ\tau}=\underline{\tau}\cdot\underline{{\sigma}}$, $(\underline{{\sigma}}x)_{j}=x_{{\sigma}(j)}$, and $(\underline{\sigma}A\underline{\tau})_{jk}=A_{\sigma(j)\tau^{-1}(k)}$ hold for all $A\in\mathbb{R}^{n\times n}$, $x\in\mathbb{R}^{n}$, $j,k=1,\ldots,n$, and all $\sigma,\tau\in S_{n}$. Now the precise connection between thermomajorization and (quasi-classical) state transfers is summarized in the following well-known result: Given any $x,y\in\mathbb{R}^{n}$ the following statements are equivalent (vom Ende and Dirr, 2022, Prop. 1).
@@
    $\sum_{i=1}^{j}x_{\tau(i)}\leq\min_{\{i=1,\ldots,n\}}\Big{(}\Big{(}\sum_{j=1}^{n}\max\Big{\{}y_{j}-\frac{y_{i}}{d_{i}}d_{j},0\Big{\}}\Big{)}+\frac{y_{i}}{d_{i}}\Big{(}\sum_{k=1}^{j}d_{\tau(k)}\Big{)}\Big{)}$
@@
[Thermomajorization for Zero Temperature] While taking the limit $T\to 0^{+}$ of $\mathrm{th}_{d,y}$ is mathematically possible—after restricting the minimum in Eq. (2) to those $i$ for which $d_{i}>0$—the resulting function does not characterize thermomajorization anymore. The reason for this is that simply taking the limit $T\to 0^{+}$ of the Gibbs vector disregards important physical aspects; indeed, taking the temperature-zero limit of the $d$-majorization conditions instead results in the correct and more restrictive notion of upper-triangular majorization Narasimhachar and Gour (2015) (cf. also p.9 in their supplemental material).
The importance of the $1$-norm conditions is that they lead to a beautiful characterization of $M_{d}(y)$ from (1), cf. (vom Ende and Dirr, 2022, Thm. 10):
@@ 3.1 Cyclic State Transfers and Subspaces in Equilibrium
Because $\mathrm{th}_{d,y}$ is concave and because $m^{\top}d\in(0,{\bf e}^{\top}d)$ by the assumptions on $m$ and $d$, Lemma A (iv) (App. A) shows that (6) can only hold if $\mathrm{th}_{d,y}$ is linear. But the latter is equivalent to $y$ being a multiple of $d$ as, by definition, the slopes of $\mathrm{th}_{d,y}$ are given by ${y_{j}}/{d_{j}}$.
@@ 3.2 Extreme Points of the Thermomajorization Polytope
Assume $\sigma\in S_{n}$ satisfies $E_{d,y}(\sigma)=y$ so $(E_{d,y}(\sigma))_{\sigma(j)}=y_{\sigma(j)}$ for all $j=1,\ldots,n$. By definition of $E_{d,y}$ this means $\mathrm{th}_{d,y}(\sum_{i=1}^{j}d_{\sigma(i)})=\mathrm{th}_{d,y}(\sum_{i=1}^{j-1}d_{\sigma(i)})+y_{\sigma(j)}$ which by induction is equivalent to $\mathrm{th}_{d,y}(\sum_{i=1}^{j}d_{\sigma(i)})=\sum_{i=1}^{j}y_{\sigma(i)}$ for all $j=1,\ldots,n-1$. But by Lemma A (App. A) this holds if and only if ${y_{\sigma(1)}}/{d_{\sigma(1)}}\geq\ldots\geq{y_{\sigma(n)}}/{d_{\sigma(n)}}$. ∎
@@ Acknowledgements.
Appendix A Basic Properties of Concave Functions